# Primal-Dual iLQR for GPU-Accelerated Learning and Control in Legged Robots

### v1

@@ Abstract
This paper introduces a novel Model Predictive Control (MPC) implementation for legged robot locomotion that leverages GPU parallelization. Our approach enables both temporal and state-space parallelization by incorporating a parallel associative scan to solve the primal-dual Karush-Kuhn-Tucker (KKT) system. In this way, the optimal control problem is solved in $\mathcal{O}(n\log{N}+m)$ complexity, instead of $\mathcal{O}(N(n+m)^{3})$, where $n$, $m$, and $N$ are the dimension of the system state, control vector, and the length of the prediction horizon. We demonstrate the advantages of this implementation over two state-of-the-art solvers (acados and crocoddyl), achieving up to a 60% improvement in runtime for Whole Body Dynamics (WB)-MPC and a 700% improvement for Single Rigid Body Dynamics (SRBD)-MPC when varying the prediction horizon length. The presented formulation scales efficiently with the problem state dimensions as well, enabling the definition of a centralized controller for up to 16 legged robots that can be computed in less than 25 ms. Furthermore, thanks to the JAX implementation, the solver supports large-scale parallelization across multiple environments, allowing the possibility of performing learning with the MPC in the loop directly in GPU.
Note: The code associated with this work can be found at: https://github.com/iit-DLSLab/mpx.
@@ I Introduction
Among the many well-known control approaches available, Model Predictive Control (MPC) has proven to be highly effective in generating and controlling complex dynamic behaviors in robotic systems, especially legged robots, as shown by [1] and [2]. At the core of an MPC is the transcription of a task we want the robot to perform into an Optimal Control Problem (OCP) and the ability to solve it fast enough to be used in a closed-loop controller.
One of the most widely used methods in robotics for solving such OCP is Differential Dynamic Programming (DDP), which gained renewed attention through the work of Todorov et al.[3], who introduced a variant known as Iterative Linear Quadratic Regulator (iLQR). iLQR discards the second-order terms of the dynamics in the Hessian, sacrificing the local quadratic convergence of DDP in favor of faster update rates. However, a key limitation of this method is its single-shooting nature, which requires a feasible initial guess as a starting point and often exhibits critical numerical issues. [4] overcame such limitation presenting a multiple shooting variant of the iLQR algorithm. More recently, [5] proposed a feasibility-driven multiple shooting approach to DDP with notable results also on real hardware [6].
@@
Another common approach to the solution of the OCP is Sequential quadratic Programming (SQP). Compared to DDP-like algorithms, SQP has been more extensively developed for general-purpose solvers due to its flexibility in handling a broader range of constraints and objectives, as well as its robustness in dealing with infeasible iterations. Solvers like OCS2 [7] and acados [8] also showed relevant results on real hardware as demonstrated in [9] and [10], thanks to the tailoring of the solver to the structure of OCP. Recent work like [11] highlights the connection between SQP and DDP, describing how the sparsity structure of the OCP can be exploited not only in DDP-like solvers but also in other solver types, such as SQP or interior point-based solvers.
@@
While the above methods have been shown to exploit the structure of the OCP, they have struggled to exploit new hardware accelerators like GPUs. Compared to CPUs, graphics cards can massively parallelize computation and are specifically designed for high-throughput linear algebra operations. Jeon et al. [12] tried to bridge the gap by developing Cusadi, a tool to convert expressions written in the well-known framework CasADI [13] to GPU. Cusadi only translates closed-form expressions into GPU-compatible code, lacking the branching and looping capabilities necessary for most solvers. [14] introduced a GPU accelerated Quadratic Programming (QP) solver, Relu-QP, leveraging the Alternating Direction Method of Multipliers (ADMM) algorithm, achieving performance superior to state-of-the-art CPU-based solutions. However, its applicability to Nonlinear Programming (NLP) is constrained by the need to pre-compute offline part of the algorithm. Lee et al. [15] implements an iLQR controller on GPU but only partially leverages its capabilities, limiting parallelization to the line search and gradient computation. In contrast, [16] focuses only on batching, solving multiple OCP s in parallel without exploiting state or temporal parallelism. [17] extends the approach of [15] by incorporating the temporal parallelization strategy introduced in [1]. However, [1] requires introducing approximations in the backward pass and performing a consensus sweep to maintain consistency. Särkkä et al. [18] showed how parallel associative scan operations can be utilized to improve the computational complexity of DDP. When parallelized on GPU, associative scans can compute the optimal control policy in $\mathcal{O}(n\log{N}+m)$ instead of the $\mathcal{O}(N(n+m)^{3})$ of the classical Riccati Recursion found in DDP, where $n$ and $m$ are the dimensions of the system state and control input, and $N$ is the prediction horizon length.
@@
In this work, we designed a SQP method that uses an associative scan-based LQR solver to solve the primal-dual Karush-Kuhn-Tucker (KKT) system efficiently. The solver avoids the offline precalculation necessary for Relu-QP [14]. Our method benefits from a multiple shooting implementation, instead of the single shooting approach presented in [18]. Our algorithm fully exploits the parallelization capabilities of GPUs both in the temporal and the state space dimensions, while maintaining, in contrast to [17], an exact backward pass. We tailored the formulation for the deployment as a receding horizon controller for legged robot locomotion and analyzed its impact on learning and control.
@@
Control:  At first, the reduction in the computational complexity of the solver may appear to have limited significance in a practical application. Some of the already described methods, like OCS2 [7] and crocoddyl [5], have been successfully used for whole-body MPC even for experiments, as shown in [19] and [2]. To ensure that such an MPC formulation can run online on real hardware, the prediction horizon length and the number of robots that can be controlled simultaneously are constrained by the time complexity of the underlying algorithm. In particular, for solvers like OCS2, acados, and crocoddyl, the complexity scales linearly with the horizon length and cubically with the state dimension. In contrast, our approach scales logarithmically with the horizon length and linearly with the state dimension. This improved scaling enables the use of more nodes at the same update rate, allowing for finer integration steps for a better approximation of the robot dynamic or, more critically, the inclusion of a Moving Horizon Estimator (MHE), such as those presented by [20] or [21], within the same optimization loop, opening the pace for ”end-to-end” MPC s capable of adapting and reacting to system changes directly using the sensor data. The reduction in complexity with respect to the state dimension enables better scalability across different robot morphologies. While other methods, such as the one presented by [22], use an approximate approach in the form of consensus ADMM to parallelize over the system state, our implementation leverages the inherent parallelism of matrix operations provided by the GPU. State space scaling becomes crucial when a centralized controller for collaborative tasks is considered. As demonstrated by [23] and [24], centralized controllers can handle complex tasks involving multiple collaborating agents. However, while these approaches achieve notable performances, the complexity of the task and the number of agents involved are constrained by their capability to solve the increasingly large OCP s at a reasonable control frequency. Our framework overcomes such barriers, as shown in Sec. IV-C.
@@
Learning:  In recent years, Reinforcement Learning (RL) has demonstrated remarkable performance, enhancing the robustness and capabilities of legged robots. RL controllers have surpassed model-based controllers in robustness against model mismatches and sensor noise. Cheng et al. [25] also demonstrated agile movements involving jumps on unstructured terrains with a real robot. However, RL policies have shown limitations in scenarios that involve crossing gaps and stepping stones, since the learning process efficiency is significantly affected due to the sparse reward signals typical of such tasks. Giftthaler et al. [26] proposed a solution to bridge the gap between learning-based and model-based controllers by incorporating MPC as a bias for the RL-policy. In contrast, works such as [27] and [28] integrated learned behaviors directly into the optimization process. Despite the noticeable results, all of the aforementioned methods that mix model-based controllers and RL suffered from a slower training process due to the dependency on CPU-based solvers. In contrast, our implementation is entirely developed in JAX [29], and it can be coupled with GPU-based simulators such as IsaacLab [30] or Mujoco XLA [31] to achieve a significant speed-up in simulation throughput.
@@ I-B Outline
This paper is organized as follows. Sec. II describes the details of the proposed solver. Sec. III presents the dynamical models used in our formulation. Sec. IV shows benefits of our approach on comparative simulations against state-of-the-art solutions. Finally, Sec. V draws the final conclusions.
@@ II-A Optimal Control Problem
$s.t.\quad$ $\boldsymbol{x}_{i+1}=h(\boldsymbol{x}_{i},\boldsymbol{u}_{i})$ (1b)
@@
where $\boldsymbol{x}_{k}$ is the state and $\boldsymbol{u}_{k}$ is the control input. $\boldsymbol{l}(\boldsymbol{x}(\cdot))$ is a quadratic cost that includes tracking and regularization terms. Finally, $\text{X}_{i}$ and $\text{U}_{i}$ are, respectively, the set of feasible state and control inputs. The remainder of this section will describe the methods we use to solve the optimization problem posed by the OCP formulation presented in (1). By leveraging both the knowledge of the problem structure and exploiting modern hardware accelerators, we are able to obtain logarithmic time complexity in horizon length $N$ as well as state $x$ and input $u$ dimensions. For simplicity, in this section, we avoid treating the inequality constraint in (1d) and (1c). These details will be treated separately in Sec. II-F. We now focus on the solution of the equality-constrain optimization problem, which only presents the initial condition (1e) and dynamics (1b) as constraints. In particular, we are going to derive a multiple-shooting approach, which, in contrast to the single-shooting one, keeps both the state and control as optimization variables. We first derive an efficient SQP algorithm that exploits the Riccati recursion to solve the equality constrained QP Sec. II-C, and then we focus on the temporal parallelization of the solver through the use of parallel associative scan, Sec. II-D.
@@ II-B Sequential Quadratic Programming
$\boldsymbol{\lambda}^{T}(\boldsymbol{x}_{i+1}-\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i}))$
where $\boldsymbol{\lambda}_{i}$ for $i=1\,\ldots\,N+1$ are the Lagrangian multipliers related to the dynamics. In an SQP method we iteratively solve the QP derived from the quadratic approximation of the Lagrangian and the linearization of the constraints at the current guess, $\boldsymbol{x}^{k}$, $\boldsymbol{u}^{k}$, where $k$ represents the last iterate. The QP is then written as:
@@
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{A}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{B}_{i}\boldsymbol{\delta u}_{i}+\boldsymbol{b}_{i}~{}~{}\text{for}\quad i=1,\dots,N$
@@
$\delta\boldsymbol{x}_{i}$ $=\boldsymbol{x}^{k+1}_{i}-\boldsymbol{x}^{k}_{i}~{}~{}~{}~{}~{}\text{for}i=0,\,\ldots,\,N+1$ (3)
$\boldsymbol{\delta u}_{i}$ $=\boldsymbol{u}^{k+1}_{i}-\boldsymbol{u}^{k}_{i}~{}~{}~{}~{}~{}\text{for}i=0,\,\ldots,\,N.$
@@ II-C Primal & dual problem solution
As shown by [32], such QP can be efficiently solved by exploiting the problem structure. By applying the Riccati Recursion, starting from the $N$ last node in the horizon and going backward in time we obtain:
@@
$\boldsymbol{P}_{i}$ $=\boldsymbol{Q}_{i}+\boldsymbol{A}_{i}^{\top}\boldsymbol{P}_{i+1}\boldsymbol{A}_{i}+\boldsymbol{K}_{i}^{\top}\boldsymbol{G}_{i}\boldsymbol{K}_{i},$
@@
$\boldsymbol{\delta u}_{i}=\boldsymbol{K}_{i}\boldsymbol{\delta x_{i}}+\boldsymbol{k}_{i}$ (5)
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{A}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{B}\boldsymbol{\delta u}_{i}+\boldsymbol{b}_{i}$

### v2

@@ Abstract
This paper introduces a novel Model Predictive Control (MPC) implementation for legged robot locomotion that leverages GPU parallelization. Our approach enables both temporal and state-space parallelization by incorporating a parallel associative scan to solve the primal-dual Karush-Kuhn-Tucker (KKT) system. In this way, the optimal control problem is solved in $\mathcal{O}(\log^{2}(n)\log{N}+\log^{2}(m))$ complexity, instead of $\mathcal{O}(N(n+m)^{3})$, where $n$, $m$, and $N$ are the dimension of the system state, control vector, and the length of the prediction horizon. We demonstrate the advantages of this implementation over two state-of-the-art solvers (acados and crocoddyl), achieving up to a 60% improvement in runtime for Whole Body Dynamics (WB)-MPC and a 700% improvement for Single Rigid Body Dynamics (SRBD)-MPC when varying the prediction horizon length. The presented formulation scales efficiently with the problem state dimensions as well, enabling the definition of a centralized controller for up to 16 legged robots that can be computed in less than 25 ms. Furthermore, thanks to the JAX implementation, the solver supports large-scale parallelization across multiple environments, allowing the possibility of performing learning with the MPC in the loop directly in GPU. The code associated with this work can be found at https://github.com/iit-DLSLab/mpx.
Index Terms:
Optimization and Optimal Control; Legged Robots; Multi-Contact Whole-Body Motion Planning and Control
@@ I Introduction
Among the many well-known control approaches available, Model Predictive Control (MPC) has proven to be highly effective in generating and controlling complex dynamic behaviors in robotic systems, especially legged robots, as shown by [Mpc] and [Perceptive_based_MPC]. At the core of an MPC is the transcription of a task we want the robot to perform into an Optimal Control Problem (OCP) and the ability to solve it fast enough to be used in a closed-loop controller.
One of the most widely used methods in robotics for solving such OCP is Differential Dynamic Programming (DDP), which gained renewed attention through the work of Todorov et al.[ilqr], who introduced a variant known as Iterative Linear Quadratic Regulator (iLQR). iLQR discards the second-order terms of the dynamics in the Hessian, sacrificing the local quadratic convergence of DDP in favor of faster update rates. However, a key limitation of this method is its single-shooting nature, which requires a feasible initial guess as a starting point and often exhibits critical numerical issues. [GNSM] overcame such limitation presenting a multiple shooting variant of the iLQR algorithm. More recently, [crocoddyl] proposed a feasibility-driven multiple shooting approach to DDP with notable results also on real hardware [talosWholeBody].
@@
Another common approach to the solution of the OCP is Sequential quadratic Programming (SQP). Compared to DDP-like algorithms, SQP has been more extensively developed for general-purpose solvers due to its flexibility in handling a broader range of constraints and objectives, as well as its robustness in dealing with infeasible iterations. Solvers like OCS2 [ocs2] and acados [acados] also showed relevant results on real hardware as demonstrated in [tamols] and [mpc_acados], thanks to the tailoring of the solver to the structure of OCP. Recent work like [sqp_lqr] highlights the connection between SQP and DDP, describing how the sparsity structure of the OCP can be exploited not only in DDP-like solvers but also in other solver types, such as SQP or interior point-based solvers.
@@
While the above methods have been shown to exploit the structure of the OCP, they have struggled to exploit new hardware accelerators like GPUs. Compared to CPUs, graphics cards can massively parallelize computation and are specifically designed for high-throughput linear algebra operations. Jeon et al. [cusadi] tried to bridge the gap by developing Cusadi, a tool to convert expressions written in the well-known framework CasADI [casadi] to GPU. Cusadi only translates closed-form expressions into GPU-compatible code, lacking the branching and looping capabilities necessary for most solvers. Bishop et al., [reluQP] introduced a GPU accelerated Quadratic Programming (QP) solver, Relu-QP, leveraging the Alternating Direction Method of Multipliers (ADMM) algorithm, achieving performance superior to state-of-the-art CPU-based solutions. However, its applicability to Nonlinear Programming (NLP) is constrained by the need to pre-compute offline part of the algorithm. Lee et al. [gpu_lqr_limited_parallelization] implements an iLQR controller on GPU but only partially leverages its capabilities, limiting parallelization to the line search and gradient computation. In contrast, [gpu_lqr_batched] focuses only on batching, solving multiple OCP s in parallel without exploiting state or temporal parallelism. [gpu_ddp] extends the approach of [gpu_lqr_limited_parallelization] by incorporating the temporal parallelization strategy introduced in [Mpc]. However, [Mpc] requires introducing approximations in the backward pass and performing a consensus sweep to maintain consistency. Frasch et al. [moritz_parallel], proposed a method to iteratively solve linear quadratic control exploiting parallel computation, while Wright et al. [partitioned_dynamics] showed how to partition dynamic programming for parallel computation. However, none of the mentioned methods achieved at the same time logarithmic time complexity and an exact formulation. On the other hand, Särkkä et al. [TemporalParallelizationLQR] showed how parallel associative scan operations can be utilized to improve the computational complexity of DDP. When parallelized on GPU, associative scans can compute the optimal control policy in $\mathcal{O}(\log^{2}(n)\log{N}+\log^{2}(m))$ instead of the $\mathcal{O}(N(n+m)^{3})$ of the classical Riccati Recursion found in DDP, where $n$ and $m$ are the dimensions of the system state and control input, and $N$ is the prediction horizon length.
@@
In this work, we designed an SQP method that uses an associative scan-based LQR solver to solve the primal-dual Karush-Kuhn-Tucker (KKT) system efficiently. The solver avoids the offline precalculation necessary for Relu-QP [reluQP]. Our method benefits from a multiple shooting implementation, instead of the single shooting approach presented in [TemporalParallelizationLQR]. Our algorithm fully exploits the parallelization capabilities of GPUs both in the temporal and the state space dimensions, while maintaining, in contrast to [gpu_ddp], an exact backward pass. We tailored the formulation for the deployment as a receding horizon controller for legged robot locomotion and analyzed its impact on learning and control.
@@
Control:  At first, the reduction in the computational complexity of the solver may appear to have limited significance in a practical application. Some of the already described methods, like OCS2 [ocs2] and crocoddyl [crocoddyl], have been successfully used for whole-body MPC even for experiments, as shown in [inverse_dynamics_mpc] and [Perceptive_based_MPC]. To ensure that such an MPC formulation can run online on real hardware, the prediction horizon length and the number of robots that can be controlled simultaneously are constrained by the time complexity of the underlying algorithm. In particular, for solvers like OCS2, acados, and crocoddyl, the complexity scales linearly with the horizon length and cubically with the state dimension. In contrast, our approach scales with the square-log of the horizon length and linearly with the state dimension. This improved scaling enables the use of more nodes at the same update rate, allowing for finer integration steps for a better approximation of the robot dynamic or, more critically, the inclusion of a Moving Horizon Estimator (MHE), such as those presented by [Invariant_moother] or [inertial_estimationlocalization], within the same optimization loop, opening the pace for ”end-to-end” MPC s capable of adapting and reacting to system changes directly using the sensor data. The reduction in complexity with respect to the state dimension enables better scalability across different robot morphologies. While other methods, such as the one presented by [AcceleratingMPC], use an approximate approach in the form of consensus ADMM to parallelize over the system state, our implementation leverages the inherent parallelism of matrix operations provided by the GPU. State space scaling becomes crucial when a centralized controller for collaborative tasks is considered. As demonstrated by [CentralizedMPC] and [LayeredMPC], centralized controllers can handle complex tasks involving multiple collaborating agents. However, while these approaches achieve notable performances, the complexity of the task and the number of agents involved are constrained by their capability to solve the increasingly large OCP s at a reasonable control frequency. Our framework overcomes such barriers, as shown in Sec. IV-C.
@@
Learning:  In recent years, Reinforcement Learning (RL) has demonstrated remarkable performance, enhancing the robustness and capabilities of legged robots. RL controllers have surpassed model-based controllers in robustness against model mismatches and sensor noise. Cheng et al. [parkour] also demonstrated agile movements involving jumps on unstructured terrains with a real robot. However, RL policies have shown limitations in scenarios that involve crossing gaps and stepping stones, since the learning process efficiency is significantly affected due to the sparse reward signals typical of such tasks. Giftthaler et al. [DTC] proposed a solution to bridge the gap between learning-based and model-based controllers by incorporating MPC as a bias for the RL-policy. In contrast, works such as [safesteps] and [learning2closethegap] integrated learned behaviors directly into the optimization process. Despite the noticeable results, all of the aforementioned methods that mix model-based controllers and RL suffered from a slower training process due to the dependency on CPU-based solvers. In contrast, our implementation is entirely developed in JAX [jaxgithub], and it can be coupled with GPU-based simulators such as IsaacLab [orbit] or Mujoco XLA [mujoco] to achieve a significant speed-up in simulation throughput.
@@ I-B Outline
This paper is organized as follows. Sec. II describes the details of the proposed solver. Sec. III presents the dynamical models used in our formulation. Sec. IV shows the benefits of our approach on comparative simulations against state-of-the-art solutions. Finally, Sec. V draws the final conclusions.
@@ II-A Optimal Control Problem
$s.t.\quad$ $\boldsymbol{x}_{i+1}=\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})$ (1b)
@@
where $\boldsymbol{x}_{i}$ is the system state and $\boldsymbol{u}_{i}$ is the control input. $\boldsymbol{l}(\boldsymbol{x_{i}},\boldsymbol{u_{i}})$ is the stage cost made of a quadratic tracking and regularization terms, while $\boldsymbol{l}(\boldsymbol{x_{N}})$ is the terminal cost, and $h(\boldsymbol{x}_{i},\boldsymbol{u}_{i})$ is the system dynamics. Finally, $\text{X}_{i}$ and $\text{U}_{i}$ are, respectively, the set of feasible system states and control inputs. The remainder of this section will describe the methods we use to solve the optimization problem posed by the OCP formulation presented in (1). By leveraging both the knowledge of the problem structure and exploiting modern hardware accelerators, we are able to obtain logarithmic time complexity in horizon length $N$ as well as square-log for the state $x$ and input $u$ dimensions. For simplicity, in this section, we avoid treating the inequality constraint in (1d) and (1c). These details will be treated separately in Sec. II-F. We now focus on the solution of the equality-constrained optimization problem, which only presents the initial condition (1e) and dynamics (1b) as constraints. In particular, we are going to derive a multiple-shooting approach, which, in contrast to the single-shooting one, keeps both the state and control as optimization variables. We first derive an efficient SQP algorithm that exploits the Riccati recursion to solve the equality constrained QP Sec. II-C, and then we focus on the temporal parallelization of the solver through the use of parallel associative scan, Sec. II-D.
@@ II-B Sequential Quadratic Programming
$\boldsymbol{{\color[rgb]{0,0,0}\lambda_{i}}}^{T}(\boldsymbol{x}_{i+1}-\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i}))$
where $\boldsymbol{\lambda}_{i}$ for $i=1\,\ldots\,N+1$ are the Lagrangian multipliers related to the dynamics. In an SQP method, we iteratively solve the QP derived from the quadratic approximation of the Lagrangian and the linearization of the constraints at the current guess, $\boldsymbol{x}^{k}$, $\boldsymbol{u}^{k}$, where $k$ represents the last iterate. The QP is then written as:
@@
$\boldsymbol{\delta x}_{i+1}=\boldsymbol{A}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{B}_{i}\boldsymbol{\delta u}_{i}+\boldsymbol{b}_{i}~~\text{for}\quad i={\color[rgb]{0,0,0}0},\dots,N$
@@
$\delta\boldsymbol{x}_{i}$ $=\boldsymbol{x}^{k+1}_{i}-\boldsymbol{x}^{k}_{i}~~~~~\text{for}\quad i=0,\,\ldots,\,N+1$ (3)
$\boldsymbol{\delta u}_{i}$ $=\boldsymbol{u}^{k+1}_{i}-\boldsymbol{u}^{k}_{i}~~~~~\text{for}\quad i=0,\,\ldots,\,N.$
@@ II-C Primal & dual problem solution
As shown by [fastGeneration], such QP can be efficiently solved by exploiting the problem structure. By applying the Riccati Recursion, starting from the $N$ last node in the horizon and going backward in time we obtain:
@@
$\boldsymbol{P}_{i}$ $=\boldsymbol{Q}_{i}+\boldsymbol{A}_{i}^{\top}\boldsymbol{P}_{i+1}\boldsymbol{A}_{i}+(\boldsymbol{S}_{i}^{\top}+\boldsymbol{A}_{i}\boldsymbol{P}_{i+1}\boldsymbol{B}_{i})\boldsymbol{K}_{i},$
@@
$\boldsymbol{\delta u}_{i}=\boldsymbol{K}_{i}\boldsymbol{\delta x_{i}}+\boldsymbol{k}_{i}\quad\boldsymbol{\delta x}_{i+1}=\boldsymbol{A}_{i}\boldsymbol{\delta x}_{i}+\boldsymbol{B}\boldsymbol{\delta u}_{i}+\boldsymbol{b}_{i}$ (5)